Nada recently set up a tax-deferred annuity to save for his retirement. He arranged to have BD 50 taken out of each of his monthly checks; it will earn 4% annual interest. She just had his 25 birthday, and his ordinary annuity comes to term when he is 65. Find the following: a. Find the future value of Nada’s annuity. (3 Marks) b. Find Nada’s total contribution to the annuity. (1 Mark) c. Find the total interest earned on the annuity.
Answer #1

Future value of the ordinary annuity = annuity [ ( 1+r)^t - 1 ] / r

r is the rate of interest

t = no of years

Future value of the ordinary annuity = 50 ( 1.04 ^40 - 1 ) / 0.04 = $4,751.28

Nada’s total contribution to the annuity = 50 *40 = $2000

total interest earned on the annuity = 4751.28 - 2000

= $ 2751.28